OpenStudy (anonymous):

Create your own factorable trinomial. Explain, in complete sentences, how the trinomial is factored and the process used to check the factors for accuracy

OpenStudy (anonymous):

\[x^{2} + 3x + 2\] First, the factors of the third term, 2, are listed. -2, -1, 1, 2 Second, pairs of factors of the third term are manipulated using addition to form a equation is is equal to the coefficient of the second term. 1 + 2 = 3 (Valid) 2 + 1 = 3 (Valid) 1 + (-2) = -1 (Invalid) 2 + (-1) = 1 (Invalid) Third, two terms are multiplied such that each term is the sum of x and one of the terms in a valid equation (from step 2). (x + 1)(x+2) = \[x^{2} + 3x + 2\] To check this, simply expand the factored side of the equation using the distributive property and verify that both sides of the equation or equivalent. \[x^{2} + 2x + x + 2\] = \[x^{2} + 3x + 2\]